# ImpedanceGPT: VLM-driven Impedance Control of Swarm of Mini-drones for Intelligent Navigation in Dynamic Environment

### v1

@@ Abstract
Swarm robotics plays a crucial role in enabling autonomous operations in dynamic and unpredictable environments. However, a major challenge remains ensuring safe and efficient navigation in environments filled with both dynamic alive (e.g., humans) and dynamic inanimate (e.g., non-living objects) obstacles. In this paper, we propose ImpedanceGPT, a novel system that combines a Vision-Language Model (VLM) with retrieval-augmented generation (RAG) to enable real-time reasoning for adaptive navigation of mini-drone swarms in complex environments.
The key innovation of ImpedanceGPT lies in the integration of VLM and RAG, which provides the drones with enhanced semantic understanding of their surroundings. This enables the system to dynamically adjust impedance control parameters in response to obstacle types and environmental conditions. Our approach not only ensures safe and precise navigation but also improves coordination between drones in the swarm.
Experimental evaluations demonstrate the effectiveness of the system. The VLM-RAG framework achieved an obstacle detection and retrieval accuracy of 80% under optimal lighting. In static environments, drones navigated dynamic inanimate obstacles at 1.4 m/s but slowed to 0.7 m/s with increased separation around humans. In dynamic environments, speed adjusted to 1.0 m/s near hard obstacles, while reducing to 0.6 m/s with higher deflection to safely avoid moving humans.
Video of ImpedanceGPT: https://youtu.be/JTdeg9bAzL4
@@
Figure 1: Illustration of the ImpedanceGPT framework for adaptive swarm navigation. The system processes a top-down view of the environment through VLM-RAG system to determine impedance parameters based on type and number of obstacles in the environment. These parameters then influence the impedance-based path planning strategy. The resulting swarm behavior demonstrates minimal deflection around rigid obstacles and greater deflection around dynamic human obstacles to ensure safe and adaptive navigation.
@@ I Introduction
Swarm robotics has emerged as a powerful solution for various autonomous applications, such as exploration, surveillance, and disaster response, where multiple robots work in concert to achieve common goals. Among the most critical challenges in swarm robotics is enabling intelligent navigation in dynamic, cluttered, and unpredictable environments. The ability to safely navigate around obstacles, whether dynamic alive (e.g., humans) or dynamic inanimate (e.g., non-living objects), while maintaining swarm coordination, remains an open problem. Traditional impedance control methods, which regulate the interaction forces between robots and obstacles, often struggle to adapt to rapidly changing environments, limiting their effectiveness in real-world scenarios.
In this paper, we introduce ImpedanceGPT, a novel system that integrates a vision-language model (VLM) with retrieval-augmented generation (RAG) – VLM-RAG – for enhanced navigation of a swarm of mini-drones in dynamic environments. The system leverages a custom database of environmental scenarios, enabling RAG to retrieve scenario-specific impedance parameters in real-time. By considering a top-down view of the environment, the system dynamically adapts to different obstacles, ensuring robust and safe navigation for the drones.
The key contributions of this work are as follows:
@@
  • An effective approach for impedance control that accounts for both dynamic alive (human) and dynamic inanimate (non-living) obstacles, enabling the drones to navigate in cluttered environments with higher safety and precision.
  • An enhancement to impedance control systems by providing real-time adaptability to changing environmental conditions and obstacle types, thus overcoming the limitations of existing methods.
This work addresses the existing challenge of adapting impedance control systems to changing and unpredictable environments, particularly with respect to the handling of both dynamic alive and dynamic inanimate obstacles. Our proposed system provides a significant step forward in swarm robotics, contributing to more intelligent, adaptive, and robust autonomous navigation in real-world scenarios.
@@ II Related Works
Swarm robotics has seen significant advancements, with many approaches aiming to solve complex tasks in dynamic environments. However, one of the major problems remains the ability of the drone swarm to maintain formation while simultaneously adapting to various obstacles and dynamic conditions of the robots [1].
Many approaches rely on a combination of global and local path planners, adding agility to control over a swarm, e.g., sampling-based methods [2] or the Artificial Potential Field (APF) algorithm [3], are utilized for drone swarm navigation, providing a way to guide agents through environments while avoiding obstacles. For example, Li [4] proposed an improved version of the APF path planning strategy for mobile robots in terms of collision avoidance and path efficiency. The use of APF-based path planning in 3D environments has been explored by Batinovic et al. [5]. Their research focused on using LiDAR sensors for obstacle detection and obstacle avoidance in UAV navigation and provided valuable insights into the integration of real-time LiDAR feedback to improve the efficiency and safety of navigation. More recently, Srivastava et al. [6] proposed a Cauchy APF method for fixed-wing UAVs that can be potentially implemented for obstacle avoidance in swarms of drones with different dynamic models.
While the above-mentioned approaches can ensure the safe navigation of swarms in areas populated with obstacles, they often neglect the issue of formation control. Several methods were proposed for local path planning and obstacle avoidance that are simultaneously able to preserve swarm patterns [7]. For example, Yu et al. [8] proposed an adaptive formation control algorithm for UAVs that includes collision avoidance and communication network switching. Their work addressed issues such as network reliability and the ability of UAVs to adapt to different environmental conditions. The idea of virtual impedance links, introduced by Hogan [9] for safe manipulation with robotic arms, was previously investigated for collaborative manipulators [10] and drone formations, e.g., Tsykunov et al. [11] proposed applying impedance links to control a swarm of micro-quadrotors. Rashad et al. [12] investigated Vision-Based Impedance Control (VBIC) for UAVs, applying it to the task of contact inspection of unknown surfaces. The topology of virtual impedance links in swarm formation was further investigated in the DandelionTouch [13] system by Fedoseev et al., addressing the interaction between drones and humans during haptic rendering of soft obstacles in virtual reality. The research highlighted the potential of using impedance-driven drone swarms to dynamically interact with moving objects and adjust impedance topology to match swarm behavior.
Figure 2: System architecture of the ImpedanceGPT.
Swarm robotics often involves multi-agent coordination in real-time, which can be challenging in cluttered environments. Darush et al. [14] presented the SwarmGear system, which utilized a morphogenetic leader drone and virtual impedance links for swarm coordination during inspection tasks. The system also demonstrated that impedance-based techniques could help to improve the stability and safety of drone swarms operating in dynamic environments. Khan et al. [15] integrated APF with dynamically adjusting impedance control for navigation in cluttered environments. The proposed SwarmPath system demonstrated the ability of impedance control to serve both pattern preservation and obstacle avoidance, however, did not take into account the varying impedance parameters of the obstacles. Zafar et al. [16] extended the application of impedance-based guidance in dynamic environments with the HetSwarm system. This research explored the cooperative navigation of heterogeneous drone swarms and emphasized the importance of task-specific guidance mechanisms for drone swarms in dynamic environments.
Recently, research has increasingly focused on integrating vision-based transformers (VBT) into robotic systems to enhance visual perception and decision-making. For example, the work by Dosovitskiy et al. [17] demonstrated the power of transformer-based models for large-scale image recognition, which has significant implications for UAV systems that rely on real-time visual data to navigate and interact with their environment. Transformers, such as those used in image processing, are crucial in enabling UAVs to handle complex visual tasks, such as object detection, scene understanding, and environment mapping. In the same context, Brohan et al. [18] introduced the Robotics Transformer (RT-1), a model designed for real-world control of robotic systems. The RT-1 model underscores the transformative potential of transformers in controlling robots across diverse tasks and environments.
Further advancements in the field are reflected in the development of Molmo-7B-D [19], a multimodal transformer model that offers state-of-the-art performance in processing diverse data types. While Molmo-7B-D is more focused on multimodal data, its core capabilities of integrating visual and other data types provide significant promise for swarm robotics, improving task execution and autonomous decision-making in cluttered and dynamic environments. Additionally, Sentence-BERT [20] and similar models have been widely adopted to allow robotic systems to better understand environmental contexts by generating efficient text embeddings. While these models are not vision-language models (VLMs) per se, they still play an important role in enabling UAVs to interpret complex multimodal data.
Our current research builds on SwarmPath by extending it into more dynamic environments, where safe and robust navigation is a critical challenge. By incorporating vision-based transformers and generative AI, we aim to enable more adaptive and real-time semantic understanding for UAV swarm navigation, allowing for greater flexibility in complex, real-world scenarios.
@@ III-A System Overview
The ImpedanceGPT framework, illustrated in Fig. 2, consists of two primary components: the VLM-RAG system for impedance parameter estimation and a high-level control system integrating an Artificial Potential Field (APF) path planner with impedance control. Inspired by the approach in [15], our implementation differs by employing a real drone as the leader.
The VLM-RAG system processes a top-down view of the environment using the Molmo-7B-D BnB 4-bit model to analyze and extract key features, including the number, type (dynamic inanimate or dynamic alive), and spatial distribution of obstacles. The extracted information is converted into a vector format and processed by the RAG system, which performs a similarity search within a custom database to retrieve optimal impedance control parameters. These parameters include virtual mass ($m$), stiffness ($k$), damping ($d$), deflection force coefficient ($F$), and separation distance ($c$) between drones. Communication between the VLM-RAG framework running on the server and the drones, as well as real-world interactions, is managed via the ROS2 framework.
@@ III-B Integrating VLM for Obstacle Identification and Spatial Analysis
The Vision-Language Model (VLM) is utilized to identify obstacles and determine their spatial arrangement. Visual inputs from a ceiling-mounted camera in the drone arena are processed to classify obstacles, quantify their number, and determine their precise locations. This enables the system to construct a semantic representation of the environment.
For efficient real-time processing, we employ the Molmo-7B-D BnB 4-bit model [19] as the VLM agent. The extracted obstacle data is then fed into the Retrieval-Augmented Generation (RAG) system, which retrieves the appropriate impedance control parameters based on the identified obstacle configuration, enhancing the swarm’s navigation and adaptability in both static and dynamic environments.
@@ III-C Custom Database for Environmental Scenarios
To build the RAG system, a custom database is developed in this work. This database contains the optimal impedance parameters, including virtual mass (m), stiffness coefficient (k), damping coefficient (d), deflection force coefficient (F), and separation distance (c) between drones, for 40 different scenarios. Each scenario represents an indoor space with varying numbers, types, and spatial arrangement of obstacles. The database is created by first simulating two and then four follower drones in a Gym PyBullet drone simulation environment, using APF as the global path planner.
@@ III-C1 Handling Dynamic Alive and Dynamic Inanimate Obstacles
In this work, humans are considered soft (dynamic alive) obstacles, while tripod stands are categorized as hard (dynamic inanimate) obstacles. The optimal impedance behavior varies based on the obstacle type:
  • Soft Obstacles: Low stiffness, moderate damping, and higher mass are required to ensure stable interactions and soft compliance with humans.
  • Hard Obstacles: High stiffness, high damping, and lower mass are needed to enable rapid response and rigid compliance.
Following these considerations, we recorded only the impedance parameters from the PyBullet simulation that met the criteria: large deflection and soft compliance for soft obstacles and small deflection and rigid compliance for hard obstacles.
Fig. 3 illustrates one of the scenarios used for data collection while Table I shows the range of these parameters for different obstacles.The lower thresholds for these parameters were determined through real-world experiments. This ensured that we established a realistic lower bound, preventing simulation experiments from going below this threshold. Additionally, this approach helped mitigate issues that arise when transferring from simulations to real-world scenarios.
Figure 3: Gym Pybullet environment for custom database collection.
TABLE I: Ranges of Optimal Parameters for Impedance Control
Parameter Dynamic Inanimate Obstacle Dynamic Alive Obstacle
Virtual Mass 1 - 1.5 3 - 7
Stiffness 7 - 10 0.1 - 0.9
Damping 3 - 5 1 - 2
Force Coefficient 0.4 - 0.7 0.2 - 0.45
Drone Separation Distance 0.2 - 0.5 0.6 - 0.9
@@ III-D Retrieval-Augmented Generation (RAG) for Impedance Parameter Generation
Retrieval-Augmented Generation (RAG) systems are typically used to enhance the response generation of Large Language Models (LLMs) by providing access to external knowledge bases. In this work, we implement a Naïve RAG approach, which utilizes a sentence transformer to create a text-embedded vector database and employs Facebook AI Similarity Search (FAISS) as the retrieval agent. FAISS is a vector search library optimized for fast nearest-neighbor searches in high-dimensional spaces using efficient indexing and search algorithms. An exact nearest-neighbor search algorithm has been used in this work.
@@
The RAG system developed in this work processes the text output of the VLM model Molmo as a query, embeds it, and then uses FAISS to retrieve the optimal impedance parameters that best match the scenario described in the query. FAISS computes similarity between the query and the embedded vectors in the database using the Euclidean distance, given by the following expression:
$d^{2}(X,Y)=\sqrt{\sum_{i=1}^{n}(X_{i}-Y_{i})^{2}}$ (1)
Where: $X=(X_{1},X_{2},...,X_{n})$ is the query embedding. $Y=(Y_{1},Y_{2},...,Y_{n})$ is the stored embedding. $n=384$ is the embedding dimension.
III-E Artificial Potential Field (APF)-Driven Global Path Planner for Leader Drone
APF is traditionally used for single-agent path planning, guiding the agent toward the goal while avoiding obstacles. In this work, we apply APF to the leader drone, which generates optimized trajectories by combining an attractive force toward the goal and a repulsive force away from obstacles. The APF equations are defined as follows [15]:
$F_{\text{total}}=F_{\text{attraction}}+F_{\text{repulsion}}$ (2)
where,
$F_{\text{attraction}}(d_{\text{g}})$ $=k_{\text{att}}\cdot d_{\text{g}},$
$F_{\text{repulsion}}(d_{\text{o}})$ =$=\begin{cases}0&\text{if }d_{\text{o}}>d_{\text{safe}}\\ k_{\text{rep}}\cdot\left(\frac{1}{d_{\text{o}}}-\frac{1}{d_{\text{safe}}}\right)&\text{if }d_{\text{o}}\leq d_{\text{safe}}\end{cases}$
where $d_{\text{g}}$ and $d_{\text{o}}$ are the distances from drone to goal and to obstacle, respectively, $k_{\text{att}}$ and $k_{\text{rep}}$ are the attraction and repulsion coefficients, respectively.
@@ III-E Artificial Potential Field (APF)-Driven Global Path Planner for Leader Drone
III-F Impedance Control for Obstacle and Collision Avoidance in Swarm Navigation
APF efficiently guides a single drone to its goal, but in a swarm, it may lead to interference and collisions. To mitigate this, we integrate APF with an impedance controller, enhancing swarm coordination and obstacle avoidance.
III-F1 Connectivity of Leader Drone with the Follower Drone
In this configuration, the follower drones positions are coupled with the leader-drone APF trajectory through a mass-spring-damper system. Thereby, creating virtual impedance links among the leader and follower drones. These impedance links provide a smooth connection between the agents and assist in preventing collisions. The links are modeled using a second-order differential equation of mass-spring-damper, which is given as:
$m\Delta\ddot{x}+d\Delta\dot{x}+k\Delta x=F_{\text{ext}}(t)$ (3)
where $\Delta x$ is the difference of current and desired drone position and $F_{\text{ext}}(t)$ is the virtual external force applied as an input from the leader. $m$ is the virtual mass of a link, $d$ is the damping coefficient of the virtual damper, and $k$ is the virtual spring constant.
III-F2 Connectivity of Follower Drones with Obstacles
In our system, the APF trajectory does not directly control the individual drones’ paths. To handle collision avoidance during local path planning, we allow the drones to detach from the APF trajectory when needed and form a temporary link with nearby obstacles, thereby pushing the drone away from it. This flexible linking mechanism ensures that the drones can avoid obstacles while rejoining the leader’s path once the obstruction is cleared:
$\Delta x_{drone,n}=k_{impF}\cdot r_{imp},$ (4)
where $r_{imp}$ is the radius of the local deflection region around the obstacle, $k_{impF}$ is the force coefficient adjusted to the drone’s average velocity, and $n$ is the drone number.

### v2

@@ Abstract
Swarm robotics plays a crucial role in enabling autonomous operations in dynamic and unpredictable environments. However, a major challenge remains ensuring safe and efficient navigation in environments shared by both dynamic alive (e.g., humans) and dynamic inanimate (e.g., non-living objects) obstacles. In this paper, we propose ImpedanceGPT, a novel system that leverages a Vision-Language Model (VLM) with Retrieval-Augmented Generation (RAG) framework to enable real-time reasoning for adaptive navigation of mini-drone swarm in complex environments. The key innovation of ImpedanceGPT lies in the merging VLM-RAG system with impedance control method, which is an active compliance strategy. This system provides drones with an enhanced semantic understanding of their surroundings and allows them to dynamically adjust impedance control parameters in response to obstacle types and environmental conditions. Our approach not only ensures safe and precise navigation but also improves coordination between drones in the swarm. Experimental evaluations demonstrate the effectiveness of our system. The VLM-RAG framework achieved an obstacle detection and retrieval accuracy of 80% under optimal lighting. In static environments, drones navigated dynamic inanimate obstacles at 1.4 m/s but slowed to 0.7 m/s with increased safety margin around humans. In dynamic environments, speed adjusted to 1.0 m/s near hard obstacles, while reducing to 0.6 m/s with higher deflection region to safely avoid moving humans.
Video of ImpedanceGPT: https://youtu.be/JTdeg9bAzL4 Github: https://github.com/Faryal-Batool/ImpedanceGPT
@@
Figure 1: ImpedanceGPT framework for adaptive swarm navigation. The system adaptively sets impedance parameters based on obstacle type and number, enabling soft compliance with alive obstacles and rigid compliance with inanimate obstacles. Impedance parameters with subscript $o$ denote links between drones and obstacles, while subscript $d$ denotes links between drones.
@@ I Introduction
Among the most critical challenges in Swarm Robotics is enabling safe and intelligent navigation in dynamic, cluttered, and unpredictable environments. The ability to safely navigate around obstacles, whether humans or dynamic non-living objects, while maintaining swarm coordination, remains an open problem. While impedance control methods offer effective and fast obstacle avoidance by shaping compliant interactions, they typically rely on manually tuned parameters and are limited in adapting to diverse obstacle types or dynamic semantic contexts. In this paper, we introduce ImpedanceGPT, a novel system that combines a Vision-Language Model (VLM) with Retrieval-Augmented Generation (RAG) method for enhanced navigation of a swarm of mini-drones in dynamic environments. The system leverages a custom database of environmental scenarios, enabling RAG to retrieve scenario-specific impedance parameters in real time. By considering a top-down view of the environment, the system dynamically adapts to different obstacles, ensuring robust and safe navigation for drones. The key contributions of this work are as follows:
  • A novel impedance control that accounts for both dynamic alive (human) and dynamic inanimate (non-living) obstacles.
@@
  • An enhancement to impedance control systems by providing real-time adaptability to changing environmental conditions and obstacle types.
@@
This work addresses the existing challenge of adapting impedance control to changing and unpredictable environments, particularly with respect to the handling of both dynamic alive and dynamic inanimate obstacles. Our proposed system provides a significant step forward in swarm robotics, contributing to more intelligent, adaptive, and robust autonomous navigation in real-world scenarios.
Figure 2: System architecture of the ImpedanceGPT. The system transmits a top-down view from a ceiling- or drone-mounted camera, along with a user request, to Molmo. Molmo identifies obstacle types, distances, and arrangements, and then passes this information to the RAG. The RAG framework searches for and retrieves impedance parameters corresponding to the scenario that best matches Molmo’s description.
@@ II Related Works
The field of safe and intelligent navigation in Swarm Robotics has significantly advanced; however, robust formation control in dynamic environments, while ensuring safe interaction within human-shared spaces, remains challenging [1]. Active compliance methods, particularly impedance control, can effectively address these limitations by dynamically adjusting impedance parameters, primarily stiffness and damping, to facilitate safe and compliant behavior [2]. Impedance control for manipulators introduced by Hogan [3] is also widely recognized for enhancing path-following accuracy and managing aerial robotic manipulation tasks. Ahmed et al. [4] introduced an Online Impedance Adaptive Control (OIAC) strategy, achieving superior UAV trajectory tracking compared to traditional PID and Model Reference Adaptive controllers. Similarly, Lippiello et al. [5] presented a Cartesian impedance control framework that enables drones to compliantly interact with external disturbances and allows manipulators to handle contact forces during aerial manipulation tasks. Further studies [6, 7, 8, 9] developed impedance control frameworks specifically tailored for aerial manipulators, where impedance behavior is applied at the manipulator’s end-effector, while the UAV itself remains rigid to maintain flight stability. Additionally, Rashad et al. [10] explored Vision-Based Impedance Control (VBIC) for UAV contact inspection tasks. However, these studies primarily focus on compliant physical interaction, with many relying on manually preselected impedance parameters or lacking validation beyond simulated environments. Intelligent impedance control was introduced in [11]. Tsykunov et al. [12] proposed the concept of virtual impedance links to achieve life-like and safe navigation of swarm of micro-quadrotors. Building on this foundation, Fedoseev et al. introduced tactile display DandelionTouch [13], demonstrating the potential for drone swarms to adapt the topology of virtual impedance links. Darush et al. [14] further demonstrated improvements in swarm stability and safety using virtual impedance links. Khan et al. [15] combined Artificial Potential Field (APF) algorithm with impedance control for navigation of drone swarm, although dynamic adjustment of impedance parameters relative to obstacle characteristics was not explored. Moreover, Zafar et al. [16] emphasized the significance of task-specific, impedance-based navigation utilizing virtual links within heterogeneous swarm configurations. Integrating impedance control with models that provide spatial and semantic understanding of the environment could further enhance the navigation capabilities. In this context, Vision-Language Models (VLMs), particularly those leveraging Vision-Based Transformers (VBTs), represent a pivotal method for augmenting visual perception and decision-making in robotic systems, significantly benefiting UAV navigation. For instance, Zhang et al. [17] proposed VLM-Social-Nav, a socially aware robot navigation framework that employs VLMs to evaluate and select navigation trajectories. Similarly, Ye et al. [18] combined VLMs with the Rapidly-exploring Random Tree (RRT) algorithm to plan paths that are not only physically feasible but also semantically meaningful. Model like Molmo-7B-O [19], capable of processing the multimodal data, has the potential to further advance autonomous decision-making in complex, dynamic environments. Additionally, Sentence-BERT [20] and similar models have been widely adopted to help robotic systems better understand environmental contexts by generating efficient text embeddings. Building upon SwarmPath [15], our current research integrates impedance control with Molmo, coupled with a RAG framework, to propose a unified system capable of predicting impedance parameters tailored to specific scenarios. This system, which can be integrated with various path planners, facilitates real-time semantic understanding for swarm navigation in complex environments, directly addressing existing research gaps where VLM-driven impedance control remains largely unexplored.
@@ III-A System Overview
The ImpedanceGPT framework, illustrated in Fig. 2, consists of two primary components: the VLM-RAG system for impedance parameter estimation and a high-level control system integrating an Artificial Potential Field (APF) path planner with impedance control. Inspired by the approach in [15], our implementation differs by employing a real drone as the leader. The VLM-RAG system processes a top-down view of the environment using the Molmo-7B-O BnB 4-bit model to analyze and extract key features, including the number, type (dynamic inanimate or dynamic alive), and spatial distribution of obstacles. The extracted information is converted into a vector format and processed by the RAG system, which performs a similarity search within a custom database to retrieve optimal impedance control parameters. These parameters ensure a collision-free path and include virtual mass ($m$), stiffness ($k$), damping ($d$), external virtual force ($F$), and separation distance ($c$) between the center of masses of the drones. Communication between the VLM-RAG framework running on the server and the drones, as well as real-world interactions, is managed via the ROS2 framework.
@@ III-B Integrating VLM for Obstacle Identification and Spatial Analysis
The VLM is employed to identify obstacles and determine their spatial arrangement. Visual inputs from a ceiling-mounted camera in the drone arena are processed to classify obstacles, count them, and determine their precise positions. This information is used to construct a semantic representation of the environment. To ensure efficient real-time processing, we utilize the Molmo-7B-O BnB 4-bit model as the VLM agent. The extracted obstacle data is then passed to the Retrieval-Augmented Generation (RAG) system, which retrieves suitable impedance control parameters based on the identified obstacle configuration. This enhances the swarm’s navigation and adaptability in both static and dynamic environments.
@@ III-C Custom Database for Environmental Scenarios
In order to build the RAG system, in this work a custom database is developed. This database contains optimal impedance parameters, including virtual mass ($m$), stiffness ($k$), damping ($d$), external virtual force ($F$), and separation distance ($c$) between the center of masses of the drones, for 40 different manually designed scenarios. Each scenario represents an indoor space with varying number, type, and spatial arrangement of obstacles. The database is created by first simulating two and then four follower drones in a Gym PyBullet drone simulation environment, using APF as the global path planner.
@@ III-C1 Handling Dynamic Alive and Dynamic Inanimate Obstacles
In this work, humans are considered as “soft” (alive) obstacles, while cylindrical obstacles are categorized as “hard” (inanimate) obstacles. The impedance behavior varies based on the obstacle type:
  • Soft (alive) obstacles: Low stiffness, moderate damping, and higher mass are required to ensure stable interactions and soft compliance with humans.
  • Hard (inanimate) obstacles: High stiffness, high damping, and lower mass are needed to enable rapid response and rigid compliance.
Following these considerations, we recorded only the optimal impedance parameters from the PyBullet simulation that met the criteria: large deflection and soft compliance for soft obstacles and small deflection and rigid compliance for hard obstacles. Table I shows the range of these parameters for different obstacles. The lower thresholds for these parameters were determined through real-world experiments.
TABLE I: Optimal Parameters for Impedance Control
Parameter
Dynamic Inanimate
Obstacle
Dynamic Alive
Obstacle
Virtual mass (kg) 1 - 1.5 3 - 7
Virtual stiffness (N/m)
7 - 10 0.1 - 0.9
Virtual damping (Ns/m)
3 - 5 1 - 2
External virtual force (N)
0.4 - 0.7 0.2 - 0.45
Separation distance (m)
0.2 - 0.5 0.6 - 0.9
@@ III-D Retrieval-Augmented Generation (RAG) for Impedance Parameter Generation
Retrieval-Augmented Generation (RAG) systems are typically used to enhance the response generation of Large Language Models (LLMs) by providing access to external knowledge bases. In this work, we implement a Naïve RAG approach, which utilizes a sentence transformer to create a text-embedded vector database and employs Facebook AI Similarity Search (FAISS) as the retrieval agent. The RAG system developed in this work processes the text output of the VLM model Molmo as a query, embeds it, and then uses FAISS to retrieve the optimal impedance parameters that best match the scenario described in the query. FAISS computes similarity between the query and the embedded vectors in the database using the Euclidean distance, given by the following expression:
$d(X,Y)=\sqrt{\sum_{i=1}^{n}(X_{i}-Y_{i})^{2}},$ (1)
where $X=(X_{1},X_{2},...,X_{n})$ is the query embedding, $Y=(Y_{1},Y_{2},...,Y_{n})$ is the stored embedding, and $n=384$ is the embedding dimension.
@@
III-E Impedance Control for Obstacle and Collision Avoidance in Drone Swarm Navigation
Impedance control is integrated with APF algorithm to enhance swarm coordination and collision avoidance in drone navigation. Follower drones maintain connectivity with the leader drone through virtual impedance links modeled as a mass-spring-damper system, described by the equation:
$m\Delta\ddot{x}+d\Delta\dot{x}+k\Delta x=F_{\text{ext}}(t),$ (2)
where $\Delta x$, $\Delta\ddot{x}$, $\Delta\ddot{x}$ are the differences between the current and desired position, velocity, and acceleration, respectively, $m$ is the virtual mass, $d$ is the virtual damping, $k$ is the virtual stiffness, and $F_{\text{ext}}(t)$ is the virtual external force from the leader drone. Additionally, drones also dynamically form temporary virtual impedance links with nearby obstacles, governed by:
$\Delta x_{\text{drone},n}=k_{impF}\cdot r_{imp},$ (3)
where $\Delta x_{\text{drone},n}$ is the difference between the obstacle and the $n_{th}$ drone; $n$ being the total number of drones, $r_{imp}$ defines the local deflection radius (0.65 m) and $k_{impF}$ adjusts the avoidance force based on drone velocity (0.45), enabling drones to temporarily diverge to avoid obstacles before rejoining trajectory of the leader drone.